OpenStudy (anonymous):

The sum of three consecutive whole numbers is a number between 70 and 100. Give three possible starting numbers for the group of three consecutive whole numbers. Solve and show all working out.

OpenStudy (mathmale):

First, express the three consecutive whole numbers algebraically. Hint: were we to begin with 20, the next consecutive whole number would be 20+1 = 21.

OpenStudy (anonymous):

So like, x + x +1 + x +2 or something?

OpenStudy (mathmale):

Yes, that's right! Simplify that result algebraically. This is your "sum." This sum is between 70 and 100. How would you express that algebraically?

OpenStudy (anonymous):

x + x + 1 +x +2 = 30?

OpenStudy (mathmale):

Why 30? We are told that the sum is between 70 and 100. Could you write an inequality expressing this fact?

OpenStudy (anonymous):

x+x+1+x+2 =70 -100?

OpenStudy (mathmale):

But that's not an inequality, is it? What does "an inequality" mean to you?

OpenStudy (anonymous):

I'm not too sure what it means....

OpenStudy (mathmale):

"The sum lies between 70 and 100" translates into the inquality \[70\le Sum \le100\]

OpenStudy (mathmale):

So, whatever your Sum is, it has to be 70 or greater or 100 or less.

OpenStudy (anonymous):

so x+x+1+x+2 = more than 70 and less than 100?

OpenStudy (mathmale):

Yes, that's right. What is the sum of x, x+1 and x+2?

OpenStudy (mathmale):

Except that you mean "equal to or more than 70 or equal to or less than 100."

OpenStudy (anonymous):

3x + 3 = equal to or more than 70 or equal to or less than 100?

OpenStudy (mathmale):

Yes, and expressing that algebraically, we get \[70\le3x+3\le100\]

OpenStudy (mathmale):

By "algebraically," I meant "symbolically."

OpenStudy (mathmale):

Remember, x will be the starting value of 3 consecutive numbers whose sum is 70 or greater , up to 100.

OpenStudy (mathmale):

What's our next step in solving for x?

OpenStudy (anonymous):

minus the three from the LHS and use the inverse operation on the RHS?

OpenStudy (mathmale):

Try this, please: just subtract 3 from the left term, the middle term and the right term (3 separate subtractions).

OpenStudy (mathmale):

@kapow...I'd be happy to help you finish this problem, but would appreciate faster responses on your part.

OpenStudy (anonymous):

Sorry about that! I will try to be more prompt with my replies!

OpenStudy (mathmale):

Sorry too. But I've waited and waited for you to subtract 3 from those 3 terms of the inequality without seeing any action on your part. Perhaps we could continue later.

OpenStudy (anonymous):

Sure! Thank you for trying to help! :)

OpenStudy (mathmale):

Next time you're not going to responding right away, please let me know. i believe I spent about ten minutes waiting for a response from you. That's time i could have spent helping other students.